OpenStudy (anonymous):

What is the value of x in the equation 4(2x - 6) =56? a. 4 b. 31/4 c.7 d. 10 please help!

OpenStudy (anonymous):

distrbute and isolate x 4(2x-8)=56 8x-32=56 +32 +32 8x=88 now just solve for x

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

oh, oops i had the wrong number 4(2x-6)=56 8x-24=56 +24 +24 8x=80

OpenStudy (anonymous):

@steffw oh XD thanks